### v1

# supplement Analyzing Defect Complex Structures of PL6 in SiC: Insights from First-Principles Screening

Abstract
The PL6 color center in 4H-SiC has attracted significant research attention owing to its remarkable spin and optical properties at ambient temperatures. Unlike many color centers that require cryogenic conditions to function, the PL6 center maintains optical activity even at room temperature or higher, benefiting from the material’s high thermal conductivity and chemical stability. To elucidate the origin of PL6’s distinctive characteristics and clarify its microscopic structure in the crystal, this study conducts a comprehensive investigation of potential defect complexes through systematic first-principles calculations. Our computational approach rigorously evaluates the electronic structures and spin-related properties of various defect configurations, enabling reasonable identification of the most probable atomic arrangement responsible for the observed characteristics of PL6. Our analysis identifies a limited set of candidate structures with -like symmetry consistent with known experimental signatures. Among these, two defect complexes emerge as the most likely PL6 candidates. The findings suggest that targeted experimental studies, such as electron spin resonance and optically detected magnetic resonance, could further clarify the nature of the PL6 center, while additional theoretical investigations on transition matrix elements and defect formation dynamics are warranted.
I Introduction
Spin defects in solid-state systems are promising candidate systems for quantum information science [1]. Among these candidates, divacancies in 4H-SiC have garnered significant attention due to their long coherence time [2, 3], spin-selective optical transitions [4], and telecom-wavelength fluorescence [5]. Photoluminescence (PL) centers, specifically PL1-8 [6, 7], have been reported in 4H-SiC. In particular, PL5 and PL6 exhibit similar optical and spin characteristics to those of PL1-4 [8, 6] but exhibit greater superior stability compared to PL1-4 under optical excitation [9, 10]. Experiment [8] and first-principles calculations [11] have unambiguously identified PL1-4 as the hh, kk, hk, and kh divacancies, but PL5-8 centers have not been fully identified.
Due to their highly similar properties, PL5 and PL6 are generally considered to be perturbed versions of the PL1-4 divacancy defects, influenced by their surrounding environment. One proposed explanation for the distinct properties of PL5 and PL6 related them to kh and kk divacancies located near stacking faults [12], while, challenged by subsequent experiments [13]. Divacancies generated via post-irradiation annealing (following electron or ion bombardment) [14] exist within a multifaceted point defect landscape in SiC, coexisting with intrinsic vacancies, antisite defects, and potential residual impurities inherent in the as-grown SiC material. Given this complex defect landscape, another plausible hypothesis emerges: PL5 and PL6 are complexes formed from a divacancy paired with another intrinsic point defect. In support of this concept, a recent study combining experimental and theoretical results proposed a new atomic model for PL5. This model identifies PL5 as a kh divacancy complexed with a nearby carbon antisite () defect [15]. Adopting this composite defect perspective – that PL6, like PL5, also involves a divacancy and another intrinsic defect – our current work aims to identify the possible atomic structures of PL6. We employ first-principles calculations within a constructed selection scheme to screen potential defect complex structures for PL6. This screening is based on zero-field splitting (ZFS), estimation of optical transition energy based on the Kohn-Sham (KS) gap, and binding energy of compounded defects. Ultimately, we identify two defect complex structures as the most probable candidates for PL6.
II Methodology
II.1 Calculation setup
The electronic structure calculations were performed using the Vienna ab initio Simulation Package (VASP) [16, 17], employing the projector augmented wave (PAW) method [18] to describe the core regions of the electronic wavefunctions with recommended PAW pseudopotentials adopted. The Revised Perdew-Burke-Ernzerhof for solids functional [19] was utilized. The plane-wave cutoff energy was set to 520 eV, with a 12122 Monkhorst-Pack k-point mesh used for the primitive cell and a single point for the expanded supercell. The convergence criterion was set to eV for the self-consistent field calculation and 0.01 eV/Å for the ionic relaxation. The ZFS matrix was calculated using a formalism that employs KS orbitals from periodic density functional calculations [20] with a home-built code. The PAW correction [21] and spin decontamination [22] were not considered here, and the accounted error was considered in the final results instead.
For the defect calculations, a cubic-like supercell was constructed to host the defect. When the defect was generated in the initial perfect cell, small random displacements were applied to all atomic coordinates, then the structure was relaxed until the force criterion was satisfied. This procedure was intended to exclude the influence of the initial configuration selection on the final obtained structure.
II.2 Determining Candidate Compounded Defects
Experimental observations show that the rhombic ZFS parameter of PL6 is almost zero [6, 8], suggesting the principal axis of PL6 is parallel to the c-axis. Consistent with this, angle-resolved measurements also indicate a -like symmetry for PL6 [13]. Based on these experimental findings and our prior considerations, hh and kk are proposed as parent models for PL6. To search for the possible defect complex structure as a variant of hh or kk, a cubic-like supercell with basis vectors (where are the basis vectors of the primitive cell) was constructed. In the first stage, a hh or kk defect was generated, with its middle placed at the center of the supercell. A cylinder with a height of 8 Å and a base radius of 7 Å centered at the defect center was chosen as the total selection range for candidates. Next, the symmetry of hh and kk was used to exclude redundant atom sites in the cylinder structure. For the remaining sites, we only consider the neutral antisite defect, i.e., the selection is conducted on the neutral complex defect consisting of hh/kk and antisite). The final selection range contains 116 cases, 56 for hh and 60 for kk. Our selection range covers all the antisites that may have a significant impact on the divacancies properties, as the spin density distribution of the defect does not uniformly decay with distance from the hh/kk center (see Fig. LABEL:spin-density in the Supplemental Material [23]).
PL6 exhibits an ground state with overall similar but distinct properties compared to PL1-4. This implies that the electron numbers of PL6 must be even, and its electronic configuration should resemble that of divacancy defects. Crucially, introducing an additional vacancy would lead to a state with , thereby definitively excluding vacancy-related configurations from consideration. Regarding antisite defects, since and act as isovalent substituents, a neutral antisite defect combined with a divacancy would alter the local structure but maintain a similar electronic environment, thus preserving the fundamental spin and optical properties. Although the charge state also satisfies the even electron number requirement, its electronic configuration significantly deviates from the divacancy case. Consequently, this study focuses exclusively on the hh/kk-type defect complexes incorporating neutral antisite defects.
II.3 Screening Methodology for the PL6 Center Identification
To systematically identify the microscopic origin of the PL6 center in SiC, we established a multi-criteria screening framework that synergistically combines spin, optical, and thermodynamic properties. The selection criteria were strategically chosen to balance computational feasibility with experimental relevance, prioritizing three key aspects: (1) Consistency with experimentally measured ZFS parameters; (2) Compatibility with photoluminescence characteristics through KS level analysis; (3) Thermodynamic stability of defect complexes. This multi-property approach enables effective discrimination among thousands of possible defect configurations while maintaining manageable computational costs.
II.3.1 ZFS Parameter Calibration and Error-Tolerant Thresholds
The ZFS matrix dominates the spin Hamiltonian of a defect center. With a proper choice of axes, it can be diagonalized, yielding two parameters that characterize the energy splitting of spin states [24]:
Note that experimentally measured values correspond to their absolute magnitudes. The ZFS parameters (, ) serve as critical fingerprints for spin defects. To address systematic errors in our DFT implementation (including PAW correction and spin decontamination errors), we performed benchmark calculations on the four well-characterized divacancies (PL1-4, i.e., hh, kk, hk, kh defects), as shown in Table 1. The calculated values are 10% systematically overestimated and the calculated values have fluctuations compared to experimental data. To ensure comprehensive coverage of potential PL6 candidates while accounting for these errors, we established conservative thresholds:
[1.365, 1.706] GHz (1.0–1.25 experimental = 1.365 GHz)
20 MHz (experimental = 0 MHz ± 20 MHz)
This error-tolerant criterion accommodates computational uncertainties and potential variations in local crystal environments.
| (GHz) | (MHz) | |||
| Divacancies | expt | calc | expt | calc |
| PL1(hh) | 1.336 | 1.506 | 0 | 6.4 |
| PL2(kk) | 1.305 | 1.455 | 0 | -4.3 |
| PL3(kh) | 1.222 | 1.395 | 82 | -61 |
| PL4(hk) | 1.334 | 1.452 | 18.6 | 29 |
II.3.2 KS Level Difference as ZPL Predictor in Defect Screening
The photoluminescence characteristics provide critical complementary signatures for defect identification. For the symmetry divacancies, the ZPL originates from the spin-allowed transition [25], corresponding to two specific occupations of KS orbitals differed by promoting an electron from the orbital to the orbital in the minority spin channel. The calculation of ZPL of PL1-4 has been detailed and tested among supercell size, k-point mesh, and functional selection [11]. However, such calculations are computationally consumption (e.g., HSE06 functional, a -only calculation with a supercell of 2400 atoms, or a calculation with -mesh for a 96-atom supercell) [11, 26].
The observed redshift in ZPL of PL6 over the hh/kk divacancies [8, 9] suggests a perturbation of the splitting between KS orbitals from an antisite defect. We establish a computationally efficient proxy for ZPL shifts through KS level analysis: the energy difference between the highest occupied () and lowest unoccupied () minority spin states systematically correlates with experimental ZPL trends in benchmarked defects (PL1-4). This relationship arises because ZPL energies and KS splitting respond proportionally to the perturbation. By requiring candidate complexes to exhibit enhanced KS splitting compared to isolated hh/kk centers, we effectively screen for structures capable of reproducing PL6’s characteristic ZPL shift.
II.3.3 Binding Energy for Stability of Complex Defects
Given that PL6 likely forms through radiation-induced defect aggregation rather than thermal equilibrium processes, we employed binding energy () rather than formation energy to assess the relative stability of candidate defects. Taking hh as an example, the binding energy is defined as [27]:
where is the formation energy of point defects. Negative values indicate exothermic complex formation, serving as a necessary (though not sufficient) condition for stability. While full thermodynamical analysis would require considering formation pathways [28] and configurational entropy [29], our screening conservatively requires to filter physically plausible complexes.
As a summary, based on experimental measurements yielding GHz and MHz for the PL6 center, we define the selection criteria for its spin properties as follows: within GHz and within MHz. To account for the observed ZPL blue shift of PL6 relative to hh and kk, we require PL6’s KS gap to exceed those of its reference hh/kk. Additionally, a negative binding energy is imposed to ensure the stability of the defect structure. Despite the intentionally permissive criteria, the geometric and electronic constraints collectively yield only a few configurations, demonstrating remarkable latent selectivity in defect screening.
III Results and Discussion
III.1 The Possible PL6 Structure
After applying the screening strategy to the 116 candidates, only 12 configurations remain, as listed in Table 2. To systematically characterize these surviving configurations, the unperturbed divacancy, and associated antisite defects are labeled by a grid-based positional encoding scheme. The numerical indices following an antisite symbol specify its position relative to the of the divacancy using grid coordinates along the -, -, and -axes, where negative displacements are denoted by overlined numbers. These coordinate axes are defined according to a gridded supercell with spacings of 1.540 Å, 0.889 Å, and 0.628 Å along each respective axis. Complete implementation details with visual examples are provided in the Supplemental Material [23].
| defect | (GHz) | (MHz) | KS gap (eV) | (eV) |
| hh | 1.506 | 6.4 | 1.117 | – |
| kk | 1.455 | -4.3 | 1.141 | – |
| 1.478 | -4.3 | 1.176 | -0.204 | |
| 1.554 | 0.3 | 1.207 | -0.083 | |
| 1.477 | 18.7 | 1.151 | -0.035 | |
| 1.464 | 1.0 | 1.149 | -0.024 | |
| 1.469 | -16.3 | 1.158 | -0.020 | |
| 1.518 | -10.0 | 1.127 | -0.014 | |
| 1.473 | 9.7 | 1.155 | -0.014 | |
| 1.516 | -11.0 | 1.129 | -0.007 | |
| 1.525 | -14.7 | 1.129 | -0.006 | |
| 1.510 | 12.6 | 1.122 | -0.004 | |
| 1.461 | 13.7 | 1.142 | -0.004 | |
| 1.459 | -13.3 | 1.142 | -0.001 |
A notable feature of the results is that all the binding energies are relatively weak, but for , with a value of eV, while all the others are at or lower than eV. This indicates that for most candidates, the attraction between the divacancy and antisite defect is weak, leading to defect complexes with limited stability. Another feature is that most of the candidates possess an whose absolute value is in the range of 10 to 20 MHz. A more stringent screening criterion on E (absolute value within 10 MHz) thus yields the following candidates: , , and . Additionally, the remaining candidates show weak KS gap enhancement relative to the original kk divacancy defect, which is consistent with the small binding energy. Among these candidates, the most enhanced KS level splitting is and . Considering all these aspects, and are the most likely PL6 candidates.
Fig. 1 shows the detailed atomic configurations of and defect complexes. Both structures comprise a kk divacancy and a antisite defect along the direction. These configurations exhibit symmetry, consistent with experimental observations indicating that PL6 is a -like center [13].
Apart from the difference in local environment between and , the dominant difference between them is the distance between the kk center and the antisite, which is 4.10 Å for and 6.00 Å for . The longer distance in leads to a weaker interaction between the kk and the antisite, which is consistent with its smaller binding energy. For these two defects, we further calculated the actual ZPL energy by using constrained DFT calculation and the 0/1- charge transition level by the PBEsol functional. While this set of computational parameters cannot reliably reproduce experimental ZPL values, it can still capture relative shifts between different configurations. This allows for a comparative analysis of trends despite the absolute discrepancy. A comparison of the results with kk is shown in Table 3. shows a larger ZPL shift of about 0.13 eV than kk, while the ZPL shift of is only about 0.03 eV. Moreover, the 0/-1 charge transition level of is closer to the CBM than kk, consistent with the experimental charge dynamics model [9]. Thus, though has a relatively small binding energy, it is the most possible PL6 candidate. Although the is closer to the kk center in , the properties of the divacancy mainly arise from the dangling bond of the C atom, which originates from the inside the divacancy. Although is farther from the kk center, it is closer to the dangling bond of the kk center, which is the reason why the impact of the antisite defect on the kk center is larger in than in .
| defect | ZPL (eV) | |
| kk | 1.024 | 0.951 |
| 1.055 | 0.962 | |
| 1.158 | 0.934 |
III.2 Potential Experimental Verification
In the above section, we identified the most promising PL6 candidates, and , through first-principles screening based on ZFS, ZPL, and binding energy. More experimental data are needed to verify the identification. In SiC, Si and C have stable isotopes with non-zero nuclear spins, namely and , with nuclear gyromagnetic ratios of -8.465 MHz/T and 10.705 MHz/T, respectively. The hyperfine interaction between nuclear spin and electron spin can be observed in ESR or ODMR spectra, providing additional signatures for defect structure identification.
The hyperfine tensor consists of an isotropic part (the Fermi contact term) and an anisotropic part , which are expressed as:
where and are the gyromagnetic ratios of the electron and nuclear spin, respectively, is the spin density, is the Dirac delta function (approximated by a smeared form here), and is the position of the nucleus [30]. The total hyperfine tensor calculations were performed using the VASP code with the PAW method and the HSE06 functional.
A key difference between and is the opposite sign of their gyromagnetic ratios. Consequently, replacing a Si atom site with a C atom leads to a significant change in the sign of the hyperfine tensor at that site. However, the sign of the hyperfine tensor is not directly observable in ESR or ODMR spectra. This stems from the fact that nuclear spin states always encompass both positive and negative components, resulting in redshifted and blueshifted fine-structure splitting pairs relative to the main peak. As the sign is difficult to determine experimentally, we adopt , where is along the c-axis (the principal axis of the tensor), as a characteristic parameter for hyperfine coupling. Furthermore, we analyze the statistical distribution of , which is independent of any specific site within the defect complex. A more detailed discussion can be found in the Supplement Material [23]. The results are presented in Fig. 2.
The statistical distribution of for is almost identical to that of kk (see Fig. LABEL:hyper_stat in Supplement Material [23]) and is omitted here for brevity. The primary difference between and kk is a redshift of approximately 0.5 MHz in the [6, 12] MHz range and a blueshift of approximately 0.7 MHz in the [12, 15] MHz range in the distribution. These shifts are attributed to the hyperfine interaction with nuclei. As outlined in Eqs. (3) and (4), the hyperfine tensor is governed by the spin density distribution, which localizes predominantly around the in the positive -direction. While the defect reverses the sign of the hyperfine interaction at its site (discussed earlier), it also induces a slight perturbation to the spin density at neighboring lattice sites. Since the component is positioned outside the region of significant spin density (as shown in Fig. LABEL:spin-density), the perturbation to the spin density is weak. This results in distribution is nearly identical to that of the isolated kk divacancy. In contrast, the distinct behavior observed for originates primarily from spatial perturbations caused by the proximal defect, rather than the sign reversal of the hyperfine interaction itself.
It is important to note that our analysis focuses on comparing the calculated hyperfine coupling of kk and , which may exhibit systematic deviations from experimental values. However, the relative shift in hyperfine coupling is experimentally observable and provides a critical means to validate our theoretical model.
III.3 Other Potential Spin Defects
| defect | spin | (GHz) | (MHz) | KS gap (eV) | (eV) | identification | |
| hh | 2 | 1.506 | 6.4 | 0.1 | 1.118 | – | PL1 |
| kk | 2 | 1.455 | -4.3 | 0.2 | 1.141 | – | PL2 |
| kh | 2 | 1.395 | -61 | 70.8 | 1.111 | – | PL3 |
| hk | 2 | 1.452 | 29 | 109.9 | 1.137 | – | PL4 |
| 0 | 0.000 | 0.0 | – | 0.555 | -4.771 | – | |
| 0 | 0.000 | 0.0 | – | 0.653 | -4.822 | – | |
| 2 | 1.545 | -26.9 | 0.0 | 1.137 | -0.964 | unknown | |
| 2 | 1.607 | -126.1 | 1.1 | 1.154 | -0.937 | unknown | |
| 2 | 1.411 | -23.3 | 177.5 | 0.951 | -0.810 | red-shifted PL | |
| 2 | 1.470 | -109.9 | 4.3 | 0.971 | -0.754 | red-shifted PL | |
| 2 | 1.287 | 144.7 | 0.8 | 0.899 | -0.387 | red-shifted PL | |
| 2 | 1.338 | 158.9 | 1.6 | 0.880 | -0.363 | red-shifted PL | |
| 2 | 0.981 | -6.9 | 0.2 | 0.603 | -0.339 | red-shifted PL | |
| 2 | 1.306 | 417.5 | 5.9 | 1.021 | -0.338 | undetectable | |
| 2 | 1.049 | 94.5 | 1.6 | 0.798 | -0.328 | red-shifted PL | |
| 2 | 1.702 | -384.4 | 7.6 | 1.038 | -0.327 | undetectable | |
| 2 | 1.347 | 459.9 | 7.0 | 1.001 | -0.321 | undetectable | |
| 2 | 1.425 | -439.8 | 11.2 | 0.664 | -0.315 | undetectable | |
| 2 | 1.769 | -433.2 | 8.5 | 1.034 | -0.309 | undetectable | |
| 2 | 1.487 | 29.7 | 0.8 | 1.040 | -0.253 | unknown | |
| 2 | 1.403 | 95.6 | 1.7 | 0.900 | -0.241 | red-shifted PL | |
| 2 | 1.544 | -27.4 | 1.0 | 1.041 | -0.238 | unknown | |
| 2 | 1.478 | -4.3 | 0.3 | 1.176 | -0.204 | PL6 | |
| 2 | 1.523 | -36.7 | 0.6 | 1.138 | -0.183 | unknown | |
| 2 | 1.460 | -43.8 | 0.2 | 1.153 | -0.159 | hk-like | |
| 2 | 1.527 | -14.8 | 179.5 | 1.111 | -0.098 | hh mixed | |
| 2 | 1.319 | -39.1 | 0.6 | 1.031 | -0.092 | unknown | |
| 2 | 1.554 | 0.3 | 0.3 | 1.207 | -0.083 | PL6 | |
| 2 | 1.387 | -32.1 | 0.7 | 1.038 | -0.082 | unknown | |
| 2 | 1.258 | 60.3 | 1.4 | 0.839 | -0.066 | red-shifted PL | |
| 2 | 1.491 | -35.4 | 1.5 | 1.098 | -0.058 | unknown | |
| 2 | 1.486 | -44.5 | 0.6 | 1.001 | -0.047 | red-shifted PL | |
| 2 | 1.484 | -23.3 | 0.2 | 0.903 | -0.041 | red-shifted PL | |
| 2 | 1.453 | 46.1 | 1.3 | 1.124 | -0.037 | hk-like | |
| 2 | 1.477 | 18.7 | 0.6 | 1.151 | -0.035 | hk-like | |
| 2 | 1.458 | -36.8 | 0.1 | 0.900 | -0.029 | red-shifted PL | |
| 2 | 1.464 | 1.0 | 179.8 | 1.149 | -0.024 | kk mixed | |
| 2 | 1.466 | -21.7 | 0.4 | 0.885 | -0.024 | red-shifted PL | |
| 2 | 1.469 | -16.3 | 0.3 | 1.158 | -0.020 | kk mixed | |
| 2 | 1.473 | 9.7 | 0.1 | 1.155 | -0.014 | kk mixed | |
| 2 | 1.518 | -10.0 | 0.1 | 1.127 | -0.014 | hh mixed | |
| 2 | 1.506 | -9.6 | 0.6 | 1.072 | -0.013 | hh mixed | |
| 2 | 1.559 | -13.2 | 0.6 | 1.109 | -0.010 | hh mixed |
In previous sections, we used experimental data to screen the possible PL6 candidates and obtain the most likely candidates. However, there are also many other defects considered in our calculation that serve as good spin defects but are not PL6 candidates. Table 4 lists all spin defects with binding energies at or below -0.01 eV, which are potentially stable, though formation dynamics were not considered. The properties of divacancies primarily originate from the C dangling bonds formed by the within the divacancy. Introducing a nearest-neighbor antisite disrupts these C dangling bonds, resulting in an ground state, as observed in the and configurations. Electron pairing in these defects enhances their stability relative to others. Besides, most -introduced defect complexes exhibit a significant reduction in crystal field splitting, leading to a lower-energy PL (less than 1.0 eV) or complete PL quenching, rendering them absent in PL spectra. Certain defects also display large (300-400 MHz), whose resonances are too widely separated in ESR or ODMR spectra for unambiguous identification. Moreover, for distant antisites, weak interactions with the central divacancy result in small perturbation, making them hard to distinguish from their parent divacancies. Notably, many defects labeled as “kh-like” exhibit spin properties similar to the hk divacancy while retaining ZPL energies near their original values. However, the principal axis of the remains approximately aligned with the c-axis, a feature resolvable through angle-resolved EPR experiments.
Among numerous stable defect complexes identified, only a limited subset demonstrates characteristics worthy of investigation as potential spin centers with distinct properties. Apart from the plausible PL6 candidates, and , eight additional stable defect complexes, highlighted in bold within Table 4, emerge as effective spin centers. Notably, the and exhibit similar structural compositions and comparable properties. Although their KS level gaps are narrower than those of PL1-4, resulting in lower ZPL energies, their and parameters remain comparable to those of PL1-4. A distinct pair, kk+ and hh+, demonstrates reduced ZPL energies alongside significantly smaller values. The most stable spin defect complexes, kk+ and hh+, exhibit ZPL energies within the PL1-4 spectral range. While kk+ shares similarities with the hk defect, it features an enhanced parameter. The hh+ complex displays an exceptionally large value and a parameter exceeding those of PL1-4 while simultaneously demonstrating superior thermodynamic stability based on binding energy analysis. Additional notable configurations include hh+(with elevated ZPL energy, increased magnitude, and higher value relative to the hh defect), and hh+(reduced ZPL energy while maintaining hk-like defect characteristics).
We have not considered exploring the transition profiles [31, 32, 33] of the PL6 candidates and other spin defects, which could serve as additional criteria. The dynamics of the formation of the defect complex could also be explored to further judge the stability of the defect complex and guide the synthesis of the defect complex in experiments.
IV Conclusion
In conclusion, we have identified the most possible PL6 candidates, and , by the first-principles screening on ZFS, ZPL, and binding energy. The antisite defect is the most probable antisite defect in the SiC crystal, and the symmetry of the defect complex is consistent with the experimental hypothesis. The experimental verification of the candidates could be conducted by the ESR or ODMR spectra, focusing on the hyperfine interaction between the nuclear spin and the electron spin. We also found eight other stable defect complexes that could serve as effective spin centers. Future work could include exploring the dynamics of the formation of the defect complex and further experimental verification of the candidates.
Acknowledgements.
The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of University of Science and Technology of China. The vaspkit code [34] was used to generate calculation files and analyze the results. Vesta [35] was used to visualize the structure.References
- Wolfowicz et al. [2021] G. Wolfowicz, F. J. Heremans, C. P. Anderson, S. Kanai, H. Seo, A. Gali, G. Galli, and D. D. Awschalom, Nat. Rev. Mater. 6, 906 (2021).
- Christle et al. [2015] D. J. Christle, A. L. Falk, P. Andrich, P. V. Klimov, J. U. Hassan, N. T. Son, E. Janzén, T. Ohshima, and D. D. Awschalom, Nat. Mater. 14, 160 (2015).
- Anderson et al. [2022] C. P. Anderson, E. O. Glen, C. Zeledon, A. Bourassa, Y. Jin, Y. Zhu, C. Vorwerk, A. L. Crook, H. Abe, J. Ul-Hassan, T. Ohshima, N. T. Son, G. Galli, and D. D. Awschalom, Sci. Adv. 8, eabm5912 (2022).
- Christle et al. [2017] D. J. Christle, P. V. Klimov, C. F. de las Casas, K. Szász, V. Ivády, V. Jokubavicius, J. U. Hassan, M. Syväjärvi, W. F. Koehl, T. Ohshima, N. T. Son, E. Janzén, Á. Gali, and D. D. Awschalom, Phys. Rev. X 7, 021046 (2017).
- Mu et al. [2021] Z. Mu, A. Rasmita, J. Yang, X. Li, and W. Gao, Adv. Quantum Technol. 4, 2100076 (2021).
- Falk et al. [2013] A. L. Falk, B. B. Buckley, G. Calusine, W. F. Koehl, V. V. Dobrovitski, A. Politi, C. A. Zorman, P. X. L. Feng, and D. D. Awschalom, Nat. Commun. 4, 1819 (2013).
- Yan et al. [2020] F.-F. Yan, A.-L. Yi, J.-F. Wang, Q. Li, P. Yu, J.-X. Zhang, A. Gali, Y. Wang, J.-S. Xu, X. Ou, C.-F. Li, and G.-C. Guo, npj Quantum Inf. 6, 38 (2020).
- Koehl et al. [2011] W. F. Koehl, B. B. Buckley, F. J. Heremans, G. Calusine, and D. D. Awschalom, Nature 479, 84 (2011).
- Wolfowicz et al. [2017] G. Wolfowicz, C. P. Anderson, A. L. Yeats, S. J. Whiteley, J. Niklas, O. G. Poluektov, F. J. Heremans, and D. D. Awschalom, Nat. Commun. 8, 1876 (2017).
- He et al. [2024] Z.-X. He, J.-Y. Zhou, Q. Li, W.-X. Lin, R.-J. Liang, J.-F. Wang, X.-L. Wen, Z.-H. Hao, W. Liu, S. Ren, H. Li, L.-X. You, R.-J. Zhang, F. Zhang, J.-S. Tang, J.-S. Xu, C.-F. Li, and G.-C. Guo, Nat. Commun. 15, 10146 (2024).
- Davidsson et al. [2018] J. Davidsson, V. Ivády, R. Armiento, N. T. Son, A. Gali, and I. A. Abrikosov, New J. Phys. 20, 023035 (2018).
- Ivády et al. [2019] V. Ivády, J. Davidsson, N. Delegan, A. L. Falk, P. V. Klimov, S. J. Whiteley, S. O. Hruszkewycz, M. V. Holt, F. J. Heremans, N. T. Son, D. D. Awschalom, I. A. Abrikosov, and A. Gali, Nat. Commun. 10, 5607 (2019).
- Son et al. [2022] N. T. Son, D. Shafizadeh, T. Ohshima, and I. G. Ivanov, J. Appl. Phys. 132, 025703 (2022).
- Luo et al. [2023] Q.-Y. Luo, Q. Li, J.-F. Wang, P.-J. Guo, W.-X. Lin, S. Zhao, Q.-C. Hu, Z.-Q. Zhu, J.-S. Xu, C.-F. Li, and G.-C. Guo, Front. Phys. 11, 10.3389/fphy.2023.1270602 (2023).
- Chen et al. [2025] Y. Chen, Q. Zhang, M. Liu, J. Liu, J. Zhou, P. Yu, S. Lin, Y. Teng, W. Yu, Y. Wang, C. Duan, F. Shi, and J. Du, (2025), under review.
- Kresse and Hafner [1993] G. Kresse and J. Hafner, Phys. Rev. B 47, 558 (1993).
- Kresse and Hafner [1994] G. Kresse and J. Hafner, Phys. Rev. B 49, 14251 (1994).
- Blöchl [1994] P. E. Blöchl, Phys. Rev. B 50, 17953 (1994).
- Perdew et al. [2008] J. P. Perdew, A. Ruzsinszky, G. I. Csonka, O. A. Vydrov, G. E. Scuseria, L. A. Constantin, X. Zhou, and K. Burke, Phys. Rev. Lett. 100, 136406 (2008).
- Rayson and Briddon [2008] M. J. Rayson and P. R. Briddon, Phys. Rev. B 77, 035119 (2008).
- Bodrog and Gali [2013] Z. Bodrog and A. Gali, J. Phys.: Condens. Matter 26, 015305 (2013).
- Biktagirov et al. [2020] T. Biktagirov, W. G. Schmidt, and U. Gerstmann, Phys. Rev. Res. 2, 022024 (2020).
- [23] See Supplement Material for further information about Possible Corresponding Defect Complex Structure of PL6 in SiC.
- Ivády et al. [2014] V. Ivády, T. Simon, J. R. Maze, I. A. Abrikosov, and A. Gali, Phys. Rev. B 90, 235205 (2014).
- Gali et al. [2010] A. Gali, A. Gällström, N. T. Son, and E. Janzén, in Mater. Sci. Forum, Vol. 645 (2010) p. 395.
- Krukau et al. [2006] A. V. Krukau, O. A. Vydrov, A. F. Izmaylov, and G. E. Scuseria, J. Chem. Phys. 125, 224106 (2006).
- Freysoldt et al. [2014] C. Freysoldt, B. Grabowski, T. Hickel, J. Neugebauer, G. Kresse, A. Janotti, and C. G. V. de Walle, Rev. Mod. Phys. 86, 253 (2014).
- Lee et al. [2021] E. M. Y. Lee, A. Yu, J. J. de Pablo, and G. Galli, Nat. Commun. 12, 6325 (2021).
- de Walle and Neugebauer [2004] C. G. V. de Walle and J. Neugebauer, J. Appl. Phys. 95, 3851 (2004).
- Blöchl [2000] P. E. Blöchl, Phys. Rev. B 62, 6158 (2000).
- Jin et al. [2021] Y. Jin, M. Govoni, G. Wolfowicz, S. E. Sullivan, F. J. Heremans, D. D. Awschalom, and G. Galli, Phys. Rev. Mater. 5, 084603 (2021).
- Razinkovas et al. [2021] L. Razinkovas, M. Maciaszek, F. Reinhard, M. W. Doherty, and A. Alkauskas, Phys. Rev. B 104, 235301 (2021).
- Quan and Duan [2025] Q. Quan and C.-K. Duan, Phys. Rev. B 111, 075156 (2025).
- Wang et al. [2021] V. Wang, N. Xu, J.-C. Liu, G. Tang, and W.-T. Geng, Comput. Phys. Commun. 267, 108033 (2021).
- Momma and Izumi [2008] K. Momma and F. Izumi, J. Appl. Crystallogr. 41, 653 (2008).

### v2

@@
-supplement Analyzing Defect Complex Structures of PL6 in SiC: Insights from First-Principles Screening
+supplement_aux Towards Identifying the PL6 Center in SiC: From First-Principles Screening to Hyperfine Validation of Competing Defect Candidates
 Abstract
-The PL6 color center in 4H-SiC has attracted significant research attention owing to its remarkable spin and optical properties at ambient temperatures.
-Unlike many color centers that require cryogenic conditions to function, the PL6 center maintains optical activity even at room temperature or higher, benefiting from the material’s high thermal conductivity and chemical stability.
-To elucidate the origin of PL6’s distinctive characteristics and clarify its microscopic structure in the crystal, this study conducts a comprehensive investigation of potential defect complexes through systematic first-principles calculations.
-Our computational approach rigorously evaluates the electronic structures and spin-related properties of various defect configurations, enabling reasonable identification of the most probable atomic arrangement responsible for the observed characteristics of PL6.
-Our analysis identifies a limited set of candidate structures with -like symmetry consistent with known experimental signatures.
-Among these, two defect complexes emerge as the most likely PL6 candidates.
-The findings suggest that targeted experimental studies, such as electron spin resonance and optically detected magnetic resonance, could further clarify the nature of the PL6 center, while additional theoretical investigations on transition matrix elements and defect formation dynamics are warranted.
+The PL6 color center in 4H-SiC, known for its excellent ambient-temperature spin and optical properties, has an unresolved microscopic origin.
+In this first-principles study, we systematically investigate potential structures to clarify its nature.
+We first rigorously examine the DV-antisite hypothesis (a divacancy paired with a carbon antisite, ), analyzing the energetic, electronic, and spin properties of various configurations.
+Two -symmetric complexes emerge as strong candidates within this framework.
+Subsequently, a critical comparison of hyperfine interaction signatures is performed between these candidates, the alternative OV model [specifically OV(hh) and OV(kk), an oxygen replacing C together with a Si vacancy], and experimental data.
+This analysis demonstrates that the OV(hh) structure more accurately reproduces PL6’s hyperfine features.
+Furthermore, re-evaluation of the proposed OV(hk) model for the related PL5 center reveals zero-field splitting parameter inconsistencies with experimental results, suggesting that PL5 and PL6 may have distinct origins.
+These findings provide crucial theoretical insights and motivate targeted experimental validation for these quantum defects.
 I Introduction
@@
 Photoluminescence (PL) centers, specifically PL1-8 [6, 7], have been reported in 4H-SiC.
-In particular, PL5 and PL6 exhibit similar optical and spin characteristics to those of PL1-4 [8, 6] but exhibit greater superior stability compared to PL1-4 under optical excitation [9, 10].
-Experiment [8] and first-principles calculations [11] have unambiguously identified PL1-4 as the hh, kk, hk, and kh divacancies, but PL5-8 centers have not been fully identified.
-Due to their highly similar properties, PL5 and PL6 are generally considered to be perturbed versions of the PL1-4 divacancy defects, influenced by their surrounding environment.
-One proposed explanation for the distinct properties of PL5 and PL6 related them to kh and kk divacancies located near stacking faults [12], while, challenged by subsequent experiments [13].
-Divacancies generated via post-irradiation annealing (following electron or ion bombardment) [14] exist within a multifaceted point defect landscape in SiC, coexisting with intrinsic vacancies, antisite defects, and potential residual impurities inherent in the as-grown SiC material.
-Given this complex defect landscape, another plausible hypothesis emerges: PL5 and PL6 are complexes formed from a divacancy paired with another intrinsic point defect.
-In support of this concept, a recent study combining experimental and theoretical results proposed a new atomic model for PL5.
-This model identifies PL5 as a kh divacancy complexed with a nearby carbon antisite () defect [15].
-Adopting this composite defect perspective – that PL6, like PL5, also involves a divacancy and another intrinsic defect – our current work aims to identify the possible atomic structures of PL6.
-We employ first-principles calculations within a constructed selection scheme to screen potential defect complex structures for PL6.
-This screening is based on zero-field splitting (ZFS), estimation of optical transition energy based on the Kohn-Sham (KS) gap, and binding energy of compounded defects.
-Ultimately, we identify two defect complex structures as the most probable candidates for PL6.
+In particular, PL5 and PL6 have similar optical and spin characteristics to those of PL1–4 [8, 6], while exhibiting superior charge stability under optical excitation [9, 10].
+Experiment [8] and first-principles calculations [11] have unambiguously identified PL1–4 as the hh, kk, hk, and kh divacancies (), where the symbols refer to the hexagonal (h) or cubic (k) crystallographic sites of the vacancy pair.
+However, the PL5-8 centers have not yet been conclusively identified.
+A schematic diagram illustrating the 4H-SiC stacking and representative vacancy complexes is shown in Fig.
+1(c).
+PL5 and PL6 exhibit properties highly similar to the PL1–4 divacancy defects and are generally considered to be divacancies perturbed by their local environment.
+Initial hypotheses proposed that PL5 and PL6 correspond to divacancies (specifically kh and kk types) located near stacking faults [12].
+However, this model was challenged by subsequent experiments [13], fueling ongoing debate and further investigation [14, 15].
+Divacancies are typically generated via post-irradiation annealing [16] of SiC samples with a multifaceted point defect landscape, including intrinsic vacancies, antisite defects, and potential residual impurities.
+Given this complex defect landscape, an alternative plausible hypothes emerges: PL5 and PL6 may be complexes formed by combining a divacancy paired with another simple point defect nearby.
+Two prominent recent proposals advance this line of thought.
+One identifies PL5 as a kh divacancy complexed with a nearby carbon antisite () [17], referred to here as the DV-antisite model.
+The other, known as the OV model, attributes both PL5 and PL6 to oxygen-related defect complexes () [18].
+Analogous to the divacancy notation, the OV configurations are differentiated by the hexagonal (h) or cubic (k) sites occupied by the and .
+Fig.
+1(c) illustrates one such configuration: (hk).
+Under the widely held assumption that PL5 and PL6 share a common origin, a critical question arises: which scenario, DV-antisite or OV, provides a more robust explanation for PL6?
+In this work, we systematically investigate the DV-antisite scenario for PL6 to identify its most probable structural origins.
+Using first-principles screening, we first analyze configurations based on zero-field splitting (ZFS), Kohn-Sham (KS) gap (used as an indicator of optical transition energy), and binding energy.
+This screening identifies two structures as promising candidates for PL6 within the DV-antisite framework.
+To further assess their validity, we compare these candidates with the leading OV(hh) structure proposed in Ref.
+[18], focusing on hyperfine interaction signatures.
+Our analysis suggests that the OV(hh) structure is the most probable origin of PL6.
+Additionally, we examine whether PL5 could be attributed to the proposed OV(hk) candidate, based on the hypothesis that PL5 and PL6 share a similar origin.
+However, inconsistencies between the calculated and measured ZFS parameter rule out the possibility that PL5 is a simple OV(hk) structure.
+The remainder of this paper is organized as follows.
+Sec.
+II outlines our comprehensive first-principles computational methodology.
+This section details the general calculation setup (Sec.
+II.1), our strategy for identifying and screening potential defect candidates for PL6 under the DV-antisite hypothesis (Sec.
+II.2 and Sec.
+II.3), and the approach used to calculate the hyperfine interactions for the subsequent validation (Sec.
+II.4).
+Sec.
+III presents our findings.
+We first discuss the results of the screening within the DV-antisite framework to identify promising candidates in Sec.
+III.1.
+This is followed by a critical validation step, comparing the hyperfine interaction signatures of DV-antisite candidates and alternative OV model structures against experimental data to determine the most probable origin of PL6 (Sec.
+III.2).
+We then extend our discussion to consider implications for the PL5 center (Sec.
+III.3).
+Finally, Sec.
+IV summarizes our conclusions.
 II Methodology
+Our investigation into the origin of PL6 commences with a systematic evaluation of the DV-antisite hypothesis, where PL6 is modeled as a divacancy paired with a nearby antisite.
+While the primary focus of the screening methodology detailed in this section is to identify promising candidates within this DV-antisite framework, the underlying first-principles calculation setup and characterization techniques are generally applicable and were consistently employed for all defect structures considered in this study, including the OV model.
+Thus, we begin by detailing the computational approaches for screening and characterizing these divacancy-antisite configurations.
 II.1 Calculation setup
@@
 This procedure was intended to exclude the influence of the initial configuration selection on the final obtained structure.
-II.2 Determining Candidate Compounded Defects
-Experimental observations show that the rhombic ZFS parameter of PL6 is almost zero [6, 8], suggesting the principal axis of PL6 is parallel to the c-axis.
+II.2 Determining Candidate Complex Defects
+To identify candidates for PL6 within the DV-antisite framework, we first consider experimental constraints.
+Experimental observations show that the rhombic ZFS parameter of PL6 is, in essence, zero within experimental uncertainties [6, 8], indicating that the principal axis of PL6 aligns with the c-axis.
 Consistent with this, angle-resolved measurements also indicate a -like symmetry for PL6 [13].
-Based on these experimental findings and our prior considerations, hh and kk are proposed as parent models for PL6.
-To search for the possible defect complex structure as a variant of hh or kk, a cubic-like supercell with basis vectors (where are the basis vectors of the primitive cell) was constructed.
-In the first stage, a hh or kk defect was generated, with its middle placed at the center of the supercell.
-A cylinder with a height of 8 Å and a base radius of 7 Å centered at the defect center was chosen as the total selection range for candidates.
-Next, the symmetry of hh and kk was used to exclude redundant atom sites in the cylinder structure.
-For the remaining sites, we only consider the neutral antisite defect, i.e., the selection is conducted on the neutral complex defect consisting of hh/kk and antisite).
-The final selection range contains 116 cases, 56 for hh and 60 for kk.
-Our selection range covers all the antisites that may have a significant impact on the divacancies properties, as the spin density distribution of the defect does not uniformly decay with distance from the hh/kk center (see Fig.
+Furthermore, crucial constraints arise from PL6’s electronic and spin properties.
+PL6 exhibits an ground state.
+This state, similar to PL1–4, implies an even electron number and an electronic configuration resembling that of divacancies.
+Consequently, our search focuses on defect complexes that preserve these features.
+Introducing an additional vacancy is excluded as it would lead to .
+While neutral antisite defects act as isovalent substituents and maintain a comparable electronic environment when combined with a diavacancy, thereby preserving the fundamental spin and optical properties, highly charged states like are excluded due to their significant deviation from the divacancy’s electronic structure.
+Therefore, this study exclusively considers neutral hh/kk-type defect complexes incorporating neutral antisite defects.
+With these constraints in mind, we proceed to search for possible complex defect structures.
+To explore variants of hh and kk configurations, we construct a cubic-like supercell with basis vectors , where are the primitive translation vectors of the primitive cell.
+In the first stage, a hh or kk defect was generated.
+Then, a cylinder was defined centered at this defect, with a height of 8 Å and a base radius of 7 Å, as the region to search for potential antisite locations.
+To reduce reduncancy, we leveraged the symmetry of the parent hh and kk divacancies to exclude symmetry-equivalent antisite locations within the cylinder.
+The final selection space consists of 116 symmetry-inequivalent antisite placements relative to the neutral divacancy: 56 for hh and 60 for kk.
+This broad configuration space is designed is capture all antisite placements with meaningful impact on the divacancy properties, considering the anisotropic spatial decay of spin density from the hh/kk center (see Fig.
 LABEL:spin-density in the Supplemental Material [26]).
-PL6 exhibits an ground state with overall similar but distinct properties compared to PL1-4.
-This implies that the electron numbers of PL6 must be even, and its electronic configuration should resemble that of divacancy defects.
-Crucially, introducing an additional vacancy would lead to a state with , thereby definitively excluding vacancy-related configurations from consideration.
-Regarding antisite defects, since and act as isovalent substituents, a neutral antisite defect combined with a divacancy would alter the local structure but maintain a similar electronic environment, thus preserving the fundamental spin and optical properties.
-Although the charge state also satisfies the even electron number requirement, its electronic configuration significantly deviates from the divacancy case.
-Consequently, this study focuses exclusively on the hh/kk-type defect complexes incorporating neutral antisite defects.
 II.3 Screening Methodology for the PL6 Center Identification
-To systematically identify the microscopic origin of the PL6 center in SiC, we established a multi-criteria screening framework that synergistically combines spin, optical, and thermodynamic properties.
-The selection criteria were strategically chosen to balance computational feasibility with experimental relevance, prioritizing three key aspects: (1) Consistency with experimentally measured ZFS parameters; (2) Compatibility with photoluminescence characteristics through KS level analysis; (3) Thermodynamic stability of defect complexes.
+To systematically identify the most plausible microscopic origin of PL6 among the candidate divacancy-antisite complexes, we established a multi-criteria screening framework integrating spin, optical, and thermodynamic properties.
+The selection criteria were designed to balance computational feasibility with experimental relevance, emphasizing three aspects: (1) consistency with experimentally measured ZFS parameters; (2) compatibility with photoluminescence characteristics through KS level analysis; (3) thermodynamic stability of defect complexes.
 This multi-property approach enables effective discrimination among thousands of possible defect configurations while maintaining manageable computational costs.
@@
 The ZFS parameters (, ) serve as critical fingerprints for spin defects.
-To address systematic errors in our DFT implementation (including PAW correction and spin decontamination errors), we performed benchmark calculations on the four well-characterized divacancies (PL1-4, i.e., hh, kk, hk, kh defects), as shown in Table 1.
-The calculated values are 10% systematically overestimated and the calculated values have fluctuations compared to experimental data.
-To ensure comprehensive coverage of potential PL6 candidates while accounting for these errors, we established conservative thresholds:
-[1.365, 1.706] GHz (1.0–1.25 experimental = 1.365 GHz)
-20 MHz (experimental = 0 MHz ± 20 MHz)
+To address systematic errors in our DFT implementation, including PAW correction and spin decontamination errors, we performed benchmark calculations on the four well-characterized divacancies (PL1–4, i.e., hh, kk, hk, kh defects), as shown in Table 1.
+The calculated values are systematically overestimated by about 10%.
+For the values, our calculations exhibit a discrepancy of roughly when compared to the experimental data.
+Notably, the value for a strictly hh or kk divacancy should be zero; however, a small nonzero value emerges due to the absence of symmetry constraints in spin-density calculations.
+To ensure comprehensive coverage of potential PL6 candidates while accounting for these errors, we adopted the following conservative thresholds:
+[1.365, 1.706] GHz ( times experimental = 1.365 GHz)
+20 MHz (allowing a MHz deviation from over the experimental value of MHz)
 This error-tolerant criterion accommodates computational uncertainties and potential variations in local crystal environments.
-| (GHz) | (MHz) | |||
-| Divacancies | expt | calc | expt | calc |
+| Divacancies | (GHz) | (MHz) | ||
+| expt | calc | expt | calc | |
 | PL1(hh) | 1.336 | 1.506 | 0 | 6.4 |
-| PL2(kk) | 1.305 | 1.455 | 0 | -4.3 |
-| PL3(kh) | 1.222 | 1.395 | 82 | -61 |
+| PL2(kk) | 1.305 | 1.455 | 0 | 4.3 |
+| PL3(kh) | 1.222 | 1.395 | 82 | 61 |
 | PL4(hk) | 1.334 | 1.452 | 18.6 | 29 |
@@
 The photoluminescence characteristics provide critical complementary signatures for defect identification.
-For the symmetry divacancies, the ZPL originates from the spin-allowed transition [25], corresponding to two specific occupations of KS orbitals differed by promoting an electron from the orbital to the orbital in the minority spin channel.
-The calculation of ZPL of PL1-4 has been detailed and tested among supercell size, k-point mesh, and functional selection [11].
-However, such calculations are computationally consumption (e.g., HSE06 functional, a -only calculation with a supercell of 2400 atoms, or a calculation with -mesh for a 96-atom supercell) [11, 26].
-The observed redshift in ZPL of PL6 over the hh/kk divacancies [8, 9] suggests a perturbation of the splitting between KS orbitals from an antisite defect.
-We establish a computationally efficient proxy for ZPL shifts through KS level analysis: the energy difference between the highest occupied () and lowest unoccupied () minority spin states systematically correlates with experimental ZPL trends in benchmarked defects (PL1-4).
+For the symmetry divacancies, the ZPL originates from the spin-allowed transition [28], corresponding to two specific occupations of KS orbitals differed by promoting an electron from the highest occupied orbital to the lowest unoccupied orbital in the minority spin channel.
+The calculation of ZPL of PL1–4 has been detailed and tested among supercell size, k-point mesh, and functional selection [11].
+However, such calculations are computationally demanding (e.g., HSE06 functional, a -only calculation with a supercell of 2400 atoms, or a calculation with -mesh for a 96-atom supercell) [11, 29].
+The observed 0.1 eV blueshift in ZPL of PL6 over the hh/kk divacancies [8, 9] suggests a perturbation of the splitting between KS orbitals from an antisite defect.
+We establish a computationally efficient proxy for ZPL shifts through KS level analysis: the energy difference between the highest occupied () and lowest unoccupied () minority spin states systematically correlates with experimental ZPL trends in benchmarked defects (PL1–4).
 This relationship arises because ZPL energies and KS splitting respond proportionally to the perturbation.
@@
 Negative values indicate exothermic complex formation, serving as a necessary (though not sufficient) condition for stability.
-While full thermodynamical analysis would require considering formation pathways [28] and configurational entropy [29], our screening conservatively requires to filter physically plausible complexes.
-As a summary, based on experimental measurements yielding GHz and MHz for the PL6 center, we define the selection criteria for its spin properties as follows: within GHz and within MHz.
-To account for the observed ZPL blue shift of PL6 relative to hh and kk, we require PL6’s KS gap to exceed those of its reference hh/kk.
-Additionally, a negative binding energy is imposed to ensure the stability of the defect structure.
-Despite the intentionally permissive criteria, the geometric and electronic constraints collectively yield only a few configurations, demonstrating remarkable latent selectivity in defect screening.
+While full thermodynamical analysis would require considering formation pathways [31] and configurational entropy [32], our screening conservatively requires to filter energetically favorable complexes.
+At this stage, by applying criteria derived from several key experimental observations for the PL6 center—namely, its ZFS parameters ( GHz and MHz) and its ZPL blueshift relative to hh/kk divacancies—along with a theoretical requirement for thermodynamic stability, we aim to identify promising candidates.
+Specifically, the applied criteria are: D within GHz, within MHz, KS gap exceeding the reference, and negative binding energy.
+These are applied to the generated divacancy-antisite configurations under the DV-antisite hypothesis.
+Despite the intentionally permissive nature of these initial criteria, the geometric and electronic constraints inherent to these specific complex defects yield only a few viable configurations, demonstrating remarkable inherent selectivity in defect screening.
+II.4 Hyperfine Interaction Calculation
+To further validate the candidates emerging from the screening process and to critically compare them against alternative models, we calculate their hyperfine interaction signatures.
+The hyperfine interation is described by a tensor comprising isotropic part (Fermi contact, ) and anisotropic () components:
+where and are the gyromagnetic ratios of the electron and nuclear spin, respectively, is the spin density, is the Dirac delta function (approximated by a smeared form here), and is the position of the nucleus [33].
+The total hyperfine tensor calculations were performed using the VASP code with the PAW method and the HSE06 functional.
+To interpret these calculations in the context of experimental data and the specific nature of our candidate defects, several considerations are pertinent.
+For the DV-antisite candidates involving a antisite, the substitution of a Si atom with a C atom means that if this antisite nucleus significantly contributes to the hyperfine spectrum, the opposite sign of the gyromagnetic ratio (compared to ) would lead to an inversion of the hyperfine coupling tensor’s sign at that particular site.
+In the case of the OV models, while the isotope possesses a nuclear spin, its extremely low natural abundance indicates that any associated hyperfine signals are too weak for detection in standard experiments without isotopic enrichment, making and the primary contributors to observable hyperfine structure.
+While the sign of the hyperfine interaction can, in principle, be determined by more complex methods like electron-nuclear double resonance (ENDOR) [34], the crucial experimental report that identified the hyperfine signatures of PL6 utilized ODMR, which resolves the magnitude of the splittings rather than the sign [35].
+Therefore, to perform the most direct and rigorous comparison with the available experimental findings, our analysis focuses on the principal component , where is the unit vector along the c-axis, coinciding with the principal axis of the ZFS tensor for the -like centers.
 III Results and Discussion
-III.1 The Possible PL6 Structure
-After applying the screening strategy to the 116 candidates, only 12 configurations remain, as listed in Table 2.
-To systematically characterize these surviving configurations, the unperturbed divacancy, and associated antisite defects are labeled by a grid-based positional encoding scheme.
+III.1 Screening Potential PL6 Structures in DV-antisite Configurations
+Applying the multi-criteria screening protocol (detailed in Sec.
+II.2) to the set of generated divacancy-antisite defect configurations, we effectively narrowed the vast configuration space.
+This process isolated only 12 structures consistent with the primary experimental constraints for PL6, as listed in Table 2.
+To systematically characterize these surviving configurations, we label the unperturbed divacancy and associated antisite defects using a grid-based positional encoding scheme.
 The numerical indices following an antisite symbol specify its position relative to the of the divacancy using grid coordinates along the -, -, and -axes, where negative displacements are denoted by overlined numbers.
-These coordinate axes are defined according to a gridded supercell with spacings of 1.540 Å, 0.889 Å, and 0.628 Å along each respective axis.
-Complete implementation details with visual examples are provided in the Supplemental Material [23].
+These coordinate axes are defined according to a gridded supercell with grid step sizes of 1.540 Å, 0.889 Å, and 0.628 Å along each respective axis.
+Complete implementation details, along with visual examples, are provided in Supplemental Material [26].
 | defect | (GHz) | (MHz) | KS gap (eV) | (eV) |
@@
 | kk | 1.455 | -4.3 | 1.141 | – |
+| OV(hh) | 1.594 | -7.2 | 1.181 | – |
+| OV(kk) | 1.523 | 2.6 | 1.158 | – |
 | 1.478 | -4.3 | 1.176 | -0.204 | |
@@
 | 1.459 | -13.3 | 1.142 | -0.001 |
-A notable feature of the results is that all the binding energies are relatively weak, but for , with a value of eV, while all the others are at or lower than eV.
-This indicates that for most candidates, the attraction between the divacancy and antisite defect is weak, leading to defect complexes with limited stability.
+A notable feature of the results is that most binding energies are small (on the order of eV or less).
+An exception is , which has a more significant binding energy of -0.204 eV.
+This reveals that for most candidates, the attraction between the divacancy and antisite defect is weak, leading to defect complexes with limited stability.
 Another feature is that most of the candidates possess an whose absolute value is in the range of 10 to 20 MHz.
-A more stringent screening criterion on E (absolute value within 10 MHz) thus yields the following candidates: , , and .
-Additionally, the remaining candidates show weak KS gap enhancement relative to the original kk divacancy defect, which is consistent with the small binding energy.
-Among these candidates, the most enhanced KS level splitting is and .
-Considering all these aspects, and are the most likely PL6 candidates.
+Considering the precise experimental value of MHz for PL6, we further refine the selection by applying a more stringent of MHz to identify candidates best matching this key experimental signature.
+This yields the following candidates: , , and .
+Additionally, the remaining candidates show weak KS gap enhancement relative to the original kk divacancy defect.
+Among these candidates, and exhibit the most enhanced KS level splitting.
+Considering all these aspects, and emerge as the most compelling candidates for the PL6 center within the configuration space of DV-antisite complexes.
 Fig.
@@
 For these two defects, we further calculated the actual ZPL energy by using constrained DFT calculation and the 0/1- charge transition level by the PBEsol functional.
-While this set of computational parameters cannot reliably reproduce experimental ZPL values, it can still capture relative shifts between different configurations.
+While this set of computational parameters cannot reliably reproduce experimental ZPL values, it should still capture the relative shifts between different configurations.
 This allows for a comparative analysis of trends despite the absolute discrepancy.
 A comparison of the results with kk is shown in Table 3.
-shows a larger ZPL shift of about 0.13 eV than kk, while the ZPL shift of is only about 0.03 eV.
-Moreover, the 0/-1 charge transition level of is closer to the CBM than kk, consistent with the experimental charge dynamics model [9].
-Thus, though has a relatively small binding energy, it is the most possible PL6 candidate.
+Notably, exhibits a significant ZPL shift of 0.13 eV relative to kk, whereas the ZPL shift for is considerably smaller, only 0.03 eV.
+Moreover, the 0/-1 charge transition level of is closer to the conduction band minimum (CBM) than kk, consistent with the experimental charge dynamics model [9].
+Thus, although has a relatively small binding energy, its ZPL shift and charge transition level characteristics establish it as a notable candidate for PL6 under the DV-antisite hypothesis.
 Although the is closer to the kk center in , the properties of the divacancy mainly arise from the dangling bond of the C atom, which originates from the inside the divacancy.
-Although is farther from the kk center, it is closer to the dangling bond of the kk center, which is the reason why the impact of the antisite defect on the kk center is larger in than in .
+Although anti-site in is farther from the kk center, it is closer to the dangling bond of the kk center, which is the reason why the impact of the antisite defect on the kk center is larger in than in .
+To summarize,, we utilized experimental data to screen potential PL6 candidates, identifying the most likely structures under the DV-antisite scenario.
+However, there are also many other defects considered in our calculation that serve as good spin defects but are not PL6 candidates.
+They are listed in Table LABEL:tab:other_spin_defects.
+Some discussion on them is also given in the Supplemental Material [26].
 | defect | ZPL (eV) | |
@@
 | 1.158 | 0.934 |
-III.2 Potential Experimental Verification
-In the above section, we identified the most promising PL6 candidates, and , through first-principles screening based on ZFS, ZPL, and binding energy.
-More experimental data are needed to verify the identification.
-In SiC, Si and C have stable isotopes with non-zero nuclear spins, namely and , with nuclear gyromagnetic ratios of -8.465 MHz/T and 10.705 MHz/T, respectively.
-The hyperfine interaction between nuclear spin and electron spin can be observed in ESR or ODMR spectra, providing additional signatures for defect structure identification.
-The hyperfine tensor consists of an isotropic part (the Fermi contact term) and an anisotropic part , which are expressed as:
-where and are the gyromagnetic ratios of the electron and nuclear spin, respectively, is the spin density, is the Dirac delta function (approximated by a smeared form here), and is the position of the nucleus [30].
-The total hyperfine tensor calculations were performed using the VASP code with the PAW method and the HSE06 functional.
-A key difference between and is the opposite sign of their gyromagnetic ratios.
-Consequently, replacing a Si atom site with a C atom leads to a significant change in the sign of the hyperfine tensor at that site.
-However, the sign of the hyperfine tensor is not directly observable in ESR or ODMR spectra.
-This stems from the fact that nuclear spin states always encompass both positive and negative components, resulting in redshifted and blueshifted fine-structure splitting pairs relative to the main peak.
-As the sign is difficult to determine experimentally, we adopt , where is along the c-axis (the principal axis of the tensor), as a characteristic parameter for hyperfine coupling.
-Furthermore, we analyze the statistical distribution of , which is independent of any specific site within the defect complex.
-A more detailed discussion can be found in the Supplement Material [23].
-The results are presented in Fig.
-2.
-The statistical distribution of for is almost identical to that of kk (see Fig.
-LABEL:hyper_stat in Supplement Material [23]) and is omitted here for brevity.
-The primary difference between and kk is a redshift of approximately 0.5 MHz in the [6, 12] MHz range and a blueshift of approximately 0.7 MHz in the [12, 15] MHz range in the distribution.
-These shifts are attributed to the hyperfine interaction with nuclei.
-As outlined in Eqs.
-(3) and (4), the hyperfine tensor is governed by the spin density distribution, which localizes predominantly around the in the positive -direction.
-While the defect reverses the sign of the hyperfine interaction at its site (discussed earlier), it also induces a slight perturbation to the spin density at neighboring lattice sites.
-Since the component is positioned outside the region of significant spin density (as shown in Fig.
-LABEL:spin-density), the perturbation to the spin density is weak.
-This results in distribution is nearly identical to that of the isolated kk divacancy.
-In contrast, the distinct behavior observed for originates primarily from spatial perturbations caused by the proximal defect, rather than the sign reversal of the hyperfine interaction itself.
-It is important to note that our analysis focuses on comparing the calculated hyperfine coupling of kk and , which may exhibit systematic deviations from experimental values.
-However, the relative shift in hyperfine coupling is experimentally observable and provides a critical means to validate our theoretical model.
-III.3 Other Potential Spin Defects
-| defect | spin | (GHz) | (MHz) | KS gap (eV) | (eV) | identification | |
-| hh | 2 | 1.506 | 6.4 | 0.1 | 1.118 | – | PL1 |
-| kk | 2 | 1.455 | -4.3 | 0.2 | 1.141 | – | PL2 |
-| kh | 2 | 1.395 | -61 | 70.8 | 1.111 | – | PL3 |
-| hk | 2 | 1.452 | 29 | 109.9 | 1.137 | – | PL4 |
-| 0 | 0.000 | 0.0 | – | 0.555 | -4.771 | – | |
-| 0 | 0.000 | 0.0 | – | 0.653 | -4.822 | – | |
-| 2 | 1.545 | -26.9 | 0.0 | 1.137 | -0.964 | unknown | |
-| 2 | 1.607 | -126.1 | 1.1 | 1.154 | -0.937 | unknown | |
-| 2 | 1.411 | -23.3 | 177.5 | 0.951 | -0.810 | red-shifted PL | |
-| 2 | 1.470 | -109.9 | 4.3 | 0.971 | -0.754 | red-shifted PL | |
-| 2 | 1.287 | 144.7 | 0.8 | 0.899 | -0.387 | red-shifted PL | |
-| 2 | 1.338 | 158.9 | 1.6 | 0.880 | -0.363 | red-shifted PL | |
-| 2 | 0.981 | -6.9 | 0.2 | 0.603 | -0.339 | red-shifted PL | |
-| 2 | 1.306 | 417.5 | 5.9 | 1.021 | -0.338 | undetectable | |
-| 2 | 1.049 | 94.5 | 1.6 | 0.798 | -0.328 | red-shifted PL | |
-| 2 | 1.702 | -384.4 | 7.6 | 1.038 | -0.327 | undetectable | |
-| 2 | 1.347 | 459.9 | 7.0 | 1.001 | -0.321 | undetectable | |
-| 2 | 1.425 | -439.8 | 11.2 | 0.664 | -0.315 | undetectable | |
-| 2 | 1.769 | -433.2 | 8.5 | 1.034 | -0.309 | undetectable | |
-| 2 | 1.487 | 29.7 | 0.8 | 1.040 | -0.253 | unknown | |
-| 2 | 1.403 | 95.6 | 1.7 | 0.900 | -0.241 | red-shifted PL | |
-| 2 | 1.544 | -27.4 | 1.0 | 1.041 | -0.238 | unknown | |
-| 2 | 1.478 | -4.3 | 0.3 | 1.176 | -0.204 | PL6 | |
-| 2 | 1.523 | -36.7 | 0.6 | 1.138 | -0.183 | unknown | |
-| 2 | 1.460 | -43.8 | 0.2 | 1.153 | -0.159 | hk-like | |
-| 2 | 1.527 | -14.8 | 179.5 | 1.111 | -0.098 | hh mixed | |
-| 2 | 1.319 | -39.1 | 0.6 | 1.031 | -0.092 | unknown | |
-| 2 | 1.554 | 0.3 | 0.3 | 1.207 | -0.083 | PL6 | |
-| 2 | 1.387 | -32.1 | 0.7 | 1.038 | -0.082 | unknown | |
-| 2 | 1.258 | 60.3 | 1.4 | 0.839 | -0.066 | red-shifted PL | |
-| 2 | 1.491 | -35.4 | 1.5 | 1.098 | -0.058 | unknown | |
-| 2 | 1.486 | -44.5 | 0.6 | 1.001 | -0.047 | red-shifted PL | |
-| 2 | 1.484 | -23.3 | 0.2 | 0.903 | -0.041 | red-shifted PL | |
-| 2 | 1.453 | 46.1 | 1.3 | 1.124 | -0.037 | hk-like | |
-| 2 | 1.477 | 18.7 | 0.6 | 1.151 | -0.035 | hk-like | |
-| 2 | 1.458 | -36.8 | 0.1 | 0.900 | -0.029 | red-shifted PL | |
-| 2 | 1.464 | 1.0 | 179.8 | 1.149 | -0.024 | kk mixed | |
-| 2 | 1.466 | -21.7 | 0.4 | 0.885 | -0.024 | red-shifted PL | |
-| 2 | 1.469 | -16.3 | 0.3 | 1.158 | -0.020 | kk mixed | |
-| 2 | 1.473 | 9.7 | 0.1 | 1.155 | -0.014 | kk mixed | |
-| 2 | 1.518 | -10.0 | 0.1 | 1.127 | -0.014 | hh mixed | |
-| 2 | 1.506 | -9.6 | 0.6 | 1.072 | -0.013 | hh mixed | |
-| 2 | 1.559 | -13.2 | 0.6 | 1.109 | -0.010 | hh mixed |
-In previous sections, we used experimental data to screen the possible PL6 candidates and obtain the most likely candidates.
-However, there are also many other defects considered in our calculation that serve as good spin defects but are not PL6 candidates.
-Table 4 lists all spin defects with binding energies at or below -0.01 eV, which are potentially stable, though formation dynamics were not considered.
-The properties of divacancies primarily originate from the C dangling bonds formed by the within the divacancy.
-Introducing a nearest-neighbor antisite disrupts these C dangling bonds, resulting in an ground state, as observed in the and configurations.
-Electron pairing in these defects enhances their stability relative to others.
-Besides, most -introduced defect complexes exhibit a significant reduction in crystal field splitting, leading to a lower-energy PL (less than 1.0 eV) or complete PL quenching, rendering them absent in PL spectra.
-Certain defects also display large (300-400 MHz), whose resonances are too widely separated in ESR or ODMR spectra for unambiguous identification.
-Moreover, for distant antisites, weak interactions with the central divacancy result in small perturbation, making them hard to distinguish from their parent divacancies.
-Notably, many defects labeled as “kh-like” exhibit spin properties similar to the hk divacancy while retaining ZPL energies near their original values.
-However, the principal axis of the remains approximately aligned with the c-axis, a feature resolvable through angle-resolved EPR experiments.
-Among numerous stable defect complexes identified, only a limited subset demonstrates characteristics worthy of investigation as potential spin centers with distinct properties.
-Apart from the plausible PL6 candidates, and , eight additional stable defect complexes, highlighted in bold within Table 4, emerge as effective spin centers.
-Notably, the and exhibit similar structural compositions and comparable properties.
-Although their KS level gaps are narrower than those of PL1-4, resulting in lower ZPL energies, their and parameters remain comparable to those of PL1-4.
-A distinct pair, kk+ and hh+, demonstrates reduced ZPL energies alongside significantly smaller values.
-The most stable spin defect complexes, kk+ and hh+, exhibit ZPL energies within the PL1-4 spectral range.
-While kk+ shares similarities with the hk defect, it features an enhanced parameter.
-The hh+ complex displays an exceptionally large value and a parameter exceeding those of PL1-4 while simultaneously demonstrating superior thermodynamic stability based on binding energy analysis.
-Additional notable configurations include hh+(with elevated ZPL energy, increased magnitude, and higher value relative to the hh defect), and hh+(reduced ZPL energy while maintaining hk-like defect characteristics).
-We have not considered exploring the transition profiles [31, 32, 33] of the PL6 candidates and other spin defects, which could serve as additional criteria.
-The dynamics of the formation of the defect complex could also be explored to further judge the stability of the defect complex and guide the synthesis of the defect complex in experiments.
+III.2 Further Validation Using the hyperfine parameter
+The previous screening identified and as key candidates for PL6 within the DV-antisite framework.
+Concurrently, the c-axis-oriented OV(hh) and OV(kk) structures, previously proposed as potential divacancy-like centers [18], also satisfy our initial screening criteria (as shown in Table 2) and thus warrant rigorous comparison.
+To further refine the identification of PL6 and critically evaluate all these leading candidates, we now analyze their predicted hyperfine interaction signatures.
+As established in our methodology (Sec.
+II.4), our analysis focuses on the principal component, , to allow for the most direct and rigorous comparison with available ODMR experimental data.
+To initially visualize and compare the overall hyperfine landscape presented by different candidate structures, we examine the statistical distribution of these calculated values for all relevant and nuclei, as depicted in Fig.
+3.
+This global representation offers an intuitive, qualitative assessment of similarities and differences in the density of hyperfine couplings across various frequency ranges.
+Fig.
+1(a) and Fig.
+1(b), which specifically depicts a kk divacancy, illustrates the typical atomic environment and key sites generally relevant for hyperfine interactions in various c-axis-oriented divacancies and related defects in 4H-SiC.
+The calculated distributions in Fig.
+3 generally exhibit characteristic regions: weakly coupled distant nuclei ([1, 4] MHz), at sites ([5, 8] MHz), and more strongly coupled at ([9, 12] MHz) and ([12, 15] MHz) sites [36], with a very strong signal from at the site around 80 MHz (omitted for clarity).
+While these overall distributions provide a valuable qualitative first look, a more quantitative and decisive comparison with available experimental data necessitates focusing on specific, well-characterized hyperfine interactions.
+Therefore, our subsequent detailed analysis and comparison with experimental findings from Ref.
+[35] will concentrate on the relative shifts of the well-characterized and sites.
+Before proceeding to a detailed comparison of specific defect models, it is crucial to establish the basis of our validation against experimental hyperfine data.
+DFT calculations of hyperfine parameters can exhibit systematic deviations from experimental absolute values due to approximations in the exchange-correlation functional, pseudopotentials, and other computational factors [37].
+However, relative shifts in hyperfine coupling strengths—such as the change observed when a nearby defect perturbs a divacancy, or differences between distinct defect configurations—are generally more robust and less susceptible to these systematic errors.
+The proven robustness of this comparative method makes it a reliable tool for validating the systems under investigation.
+Therefore, our primary validation strategy will focus on comparing these calculated relative shifts with experimentally observed trends, rather than on achieving precise agreement in absolute coupling constants.
+This approach allows for a more reliable assessment of which candidate structure best reproduces the experimentally characterized hyperfine signatures of PL6.
+With this validation approach in mind, initial examination reveals that the distribution for the complex is nearly indistinguishable from that of the unperturbed kk divacancy (see Fig.
+LABEL:hyper_stat in the Supplemental Material [26](a)), making it an improbable source for the distinct hyperfine signature of PL6.
+More significantly, the candidate, while initially promising based on other metrics, shows a notable discrepancy when its hyperfine predictions are compared to experiment.
+Specifically, our calculations for this complex predict a redshift of the peak (0.5 MHz) but a blueshift of the peak (0.7 MHz) relative to the kk divacancy.
+This pattern is inconsistent with the experimental findings for PL6, which report redshifts for both the (0.4 MHz) and (0.7 MHz) sites when compared to the kk reference [35].
+Turning to the OV candidates, the OV(kk) complex can also be dismissed.
+Our calculations predict it would induce blueshifts at both sites(0.5 MHz for and 0.4 MHz for ) relative to the kk divacancy, again conflicting with the experimental shifts of PL6.
+In striking contrast, the OV(hh) model exhibits excellent agreement with experimental trends.
+Our calculations for OV(hh) predict redshifts of 0.4 MHz for and 0.6 MHz for sites relative to the kk divacancy.
+Furthermore, when comparing to the hh divacancy as another reference, PL6 experimentally exhibits blueshifted and hyperfine couplings; the OV(hh) model successfully reproduces this, yielding calculated blueshifts of 0.5 MHz() and 0.2 MHz ().
+This consistent agreement in relative hyperfine shifts across multiple reference points and specific sites is compelling.
+Thus, the hyperfine analysis, grounded in the comparison of these robust relative changes, provides strong evidence identifying OV(hh) as the PL6 center.
+III.3 Discussion on the PL5 center
+Our preceding analysis, particularly the hyperfine validation comparing potential DV-antisite structures against OV models, strongly indicates that OV(hh) is the PL6 center.
+An OV(hk) origin for the PL5 center was proposed in Ref.
+[18], based on its calculated ZFS value, ZPL, and formation energy.
+However, for a basal defect like hk- (and by extension, OV(hk)), the nonzero ZFS value is also a critical parameter for identification, which has not been explicitly discussed for OV(hk) previously.
+We therefore calculated the ZFS value for OV(hk) and found its magnitude to be 96.9 MHz.
+Similarly, an value of 150 MHz has been reported by another independent calculation [38].
+Compared to the experimental value for PL5, which is approximately only 16 MHz [6, 39], the significantly larger calculated values (96.9 MHz and 150 MHz) represent a pronounced discrepancy that is unlikely to stem from computational inaccuracies alone.
+Therefore, our findings do not support the OV(hk) configuration as the origin of the PL5 center.
+These findings have notable implications, as PL5 and PL6 are often assumed to share a common origin due to their closely matched ZPL energies [6, 39, 8], stability in charge states [9], and qualitatively similar relative concentrations [17] observed experimentally.
+However, our results indicate a divergence: while OV(hh) emerges as the most probable candidate for PL6, the PL5 center cannot be straightforwardly attributed to a simple OV(hk) structure.
+However, it remains possible that PL5 originates from a more intricate configuration beyond the model proposed in Ref.
+[17].
+The potential distinction in origin between PL5 and PL6 can be considered in at least two ways:
+Both PL5 and PL6 are fundamentally related to oxygen-vacancy, but PL5 involves a more complex structural variation than the simple OV(hk), thereby maintaining a ‘common family’ association.
+PL5 and PL6 arise from fundamentally different sources: PL6 is OV(hh)-based, whereas PL5 may originate from an entirely different defect type, such as a divacancy-antisite complex ().
+Distinguishing between these two scenarios is essential.
+Future experimental investigations, including controlled oxygen doping and subsequent analysis of PL5 and PL6 concentrations and properties, could yield valuable insights.
+Additionally, theoretical studies should extend the search to encompass more intricate oxygen-related defect configurations beyond simple structures for PL5, complemented by experiment explorations.
 IV Conclusion
-In conclusion, we have identified the most possible PL6 candidates, and , by the first-principles screening on ZFS, ZPL, and binding energy.
-The antisite defect is the most probable antisite defect in the SiC crystal, and the symmetry of the defect complex is consistent with the experimental hypothesis.
-The experimental verification of the candidates could be conducted by the ESR or ODMR spectra, focusing on the hyperfine interaction between the nuclear spin and the electron spin.
-We also found eight other stable defect complexes that could serve as effective spin centers.
-Future work could include exploring the dynamics of the formation of the defect complex and further experimental verification of the candidates.
+In conclusion, we have performed a systematic first-principles investigation towards elucidating the microscopic origin of the PL6 center in 4H-SiC.
+Our study initially examined the DV-antisite hypothesis, screening divacancy-antisite () configurations based on zero-field splitting (ZFS), Kohn-Sham (KS) gap, and binding energy.
+This analysis identified two promising structures, specifically and , as leading candidates within this framework, both exhibiting the requisite symmetry consistent with experimental observations.
+To further validate these candidates, we performed a critical comparison of their predicted hyperfine interaction signatures of the most promising DV-antisite candidates with experimental data for PL6 and with those calculated for the alternative OV models.
+This comparative analysis demonstrated that the OV(hh) structure accurately reproduces the experimental hyperfine signatures, identifying it as the origin of PL6.
+Additionally, our re-evaluation of the proposed OV(hk) structure for PL5 identified significant discrepancies between its calculated ZFS parameter and experimental observations, raising doubt about its viability as the PL5 center.
+This indicates that PL5 and PL6 have distinct microscopic origins.
+These findings underscore the necessity of integrating multiple theoretical metrics with diverse experimental signatures to achieve reliable defect identification.
+Future targeted ESR/ODMR experiments, specifically designed to probe the distinct hyperfine signatures of OV(hh) versus complexes, will be crucial for definitively confirming the PL6 structure.
+Furthermore, our broader screening has identified additional divacancy-antisite complexes that, while not corresponding to PL6, may represent other stable spin defects in SiC worthy of future investigation.
 Acknowledgements.
+This work was supported by Innovation Program for Quantum Science and Technology (Grant Nos.
+2021ZD0302200 and 2021ZD0303204), the National Natural Science Foundation of China (Grant Nos.
+12474242 and 12274396), the Anhui Initiative in Quantum Information Technologies (Grant No.
+AHY050000), the CAS Project for Young Scientists in Basic Research, the Fundamental Research Funds for the Central Universities, and the China Postdoctoral Science Foundation (Grant Nos.
+2021M703110 and 2022T150631).
 The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of University of Science and Technology of China.
@@
 - Mu et al. [2021] Z. Mu, A. Rasmita, J. Yang, X. Li, and W. Gao, Adv. Quantum Technol. 4, 2100076 (2021).
-- Falk et al. [2013] A. L. Falk, B. B. Buckley, G. Calusine, W. F. Koehl, V. V. Dobrovitski, A. Politi, C. A. Zorman, P. X. L. Feng, and D. D. Awschalom, Nat. Commun. 4, 1819 (2013).
+- Falk et al. [2013] A. L. Falk, B. B. Buckley, G. Calusine, W. F. Koehl, V. V. Dobrovitski, A. Politi, C. A. Zorman, P. X.-L. Feng, and D. D. Awschalom, Nat. Commun. 4, 1819 (2013).
 - Yan et al. [2020] F.-F. Yan, A.-L. Yi, J.-F. Wang, Q. Li, P. Yu, J.-X. Zhang, A. Gali, Y. Wang, J.-S. Xu, X. Ou, C.-F. Li, and G.-C. Guo, npj Quantum Inf. 6, 38 (2020).
@@
 - Son et al. [2022] N. T. Son, D. Shafizadeh, T. Ohshima, and I. G. Ivanov, J. Appl. Phys. 132, 025703 (2022).
-- Luo et al. [2023] Q.-Y. Luo, Q. Li, J.-F. Wang, P.-J. Guo, W.-X. Lin, S. Zhao, Q.-C. Hu, Z.-Q. Zhu, J.-S. Xu, C.-F. Li, and G.-C. Guo, Front. Phys. 11, 10.3389/fphy.2023.1270602 (2023).
-- Chen et al. [2025] Y. Chen, Q. Zhang, M. Liu, J. Liu, J. Zhou, P. Yu, S. Lin, Y. Teng, W. Yu, Y. Wang, C. Duan, F. Shi, and J. Du, (2025), under review.
+- Shafizadeh et al. [2025] D. Shafizadeh, N. T. Son, I. A. Abrikosov, and I. G. Ivanov, Phys. Rev. B 111, 165201 (2025).
+- Li et al. [2025] J. Li, H. Qi, F. Zhou, Y. Song, N. Xu, B. Hong, H. Chen, Y. Dong, and X. Wang, arXiv:2503.22270 (2025).
+- Luo et al. [2023] Q.-Y. Luo, Q. Li, J.-F. Wang, P.-J. Guo, W.-X. Lin, S. Zhao, Q.-C. Hu, Z.-Q. Zhu, J.-S. Xu, C.-F. Li, and G.-C. Guo, Front. Phys. 11, 1270602 (2023).
+- Chen et al. [2025] Y. Chen, Q. Zhang, M. Liu, J. Liu, J. Zhou, P. Yu, S. Lin, Y. Teng, W. Yu, Y. Wang, et al., arXiv:2504.07558 (2025).
+- Bai et al. [2025] R. Bai, D. Liu, S. Chen, and Y.-N. Wu, Phys. Rev. B 111, 014106 (2025).
 - Kresse and Hafner [1993] G. Kresse and J. Hafner, Phys. Rev. B 47, 558 (1993).
@@
 - Blöchl [2000] P. E. Blöchl, Phys. Rev. B 62, 6158 (2000).
-- Jin et al. [2021] Y. Jin, M. Govoni, G. Wolfowicz, S. E. Sullivan, F. J. Heremans, D. D. Awschalom, and G. Galli, Phys. Rev. Mater. 5, 084603 (2021).
-- Razinkovas et al. [2021] L. Razinkovas, M. Maciaszek, F. Reinhard, M. W. Doherty, and A. Alkauskas, Phys. Rev. B 104, 235301 (2021).
-- Quan and Duan [2025] Q. Quan and C.-K. Duan, Phys. Rev. B 111, 075156 (2025).
+- Murzakhanov et al. [2021] F. F. Murzakhanov, B. V. Yavkin, G. V. Mamin, S. B. Orlinskii, H. J. von Bardeleben, T. Biktagirov, U. Gerstmann, and V. A. Soltamov, Phys. Rev. B 103, 245203 (2021).
+- Falk et al. [2015] A. L. Falk, P. V. Klimov, V. Ivády, K. Szász, D. J. Christle, W. F. Koehl, A. Gali, and D. D. Awschalom, Phys. Rev. Lett. 114, 247603 (2015).
+- Son et al. [2006] N. T. Son, P. Carlsson, J. ul Hassan, E. Janzén, T. Umeda, J. Isoya, A. Gali, M. Bockstedte, N. Morishita, T. Ohshima, and H. Itoh, Phys. Rev. Lett. 96, 055501 (2006).
+- Ivády et al. [2018] V. Ivády, I. A. Abrikosov, and A. Gali, npj Comput. Mater. 4, 76 (2018).
+- Kobayashi et al. [2023] T. Kobayashi, T. Shimura, and H. Watanabe, J. Appl. Phys. 134, 145701 (2023).
+- Li et al. [2021] Q. Li, J.-F. Wang, F.-F. Yan, J.-Y. Zhou, H.-F. Wang, H. Liu, L.-P. Guo, X. Zhou, A. Gali, Z.-H. Liu, et al., Natl. Sci. Rev. 9, nwab122 (2021).
 - Wang et al. [2021] V. Wang, N. Xu, J.-C. Liu, G. Tang, and W.-T. Geng, Comput. Phys. Commun. 267, 108033 (2021).